# Demystifying Workload Imbalances in Large Transformer Model Training over Variable-length Sequences

### v1

@@ 5.3 Subgraphs
To tackle the challenge of representing diverse heterogeneous strategies and their complex optimization-propagation interactions, we introduce subgraphs— subsets of the complete computation graph. The key observation is that different strategies only have unique patterns in some specific operators, while most user-defined computational operators remain unchanged. Therefore, subgraphs are developed to encapsulate these unique operators from different strategies for efficient representation, management and scheduling.
Figure 8: Scheduling subgraphs with pipeline parallel (PP). In the first PP stage, Pull Parameter subgraphs are scheduled immediately after the last operators requiring the parameter, and Push Gradient subgraphs follow right after the operators that produce or accumulate the gradient. In subsequent PP stages, Pull Parameter and Push Gradient are grouped at the beginning and end, respectively, to utilize inherent PP bubbles. This scheduling scheme effectively hides communication latency within the Pull Parameter and Push Gradient subgraphs.
@@
Subgraphs instantiation. As illustrated at the bottom of Figure 2(c), users start by defining various computational operators to describe the training model. We then identify all potential points in the computational graph where different heterogeneous strategies might intervene and preserve several empty subgraphs (e.g., between transformer modules for potential pipeline parallel, before and after matrix multiplication for potential tensor parallel). Given a strategy $S$, the user-defined operators are directly integrated into a new, strategy-specific graph. We then determine how to instantiate all predefined subgraphs with appropriate operators. For example, as depicted at the top of Figure 2(c), a data transfer operator and a strategy-specific LNK operator are added to instantiate a Pull Parameter subgraph. This instantiation process ultimately forms a complete graph for a specific $S$, enabling different GPUs to extract and execute their respective operators.
Subgraphs scheduling. During execution, we optimize overlap by controlling the execution order of subgraphs and computational operators. Each subgraph can be treated as a cohesive unit and integrated into the scheduler as a fused operator. As illustrated in Figure 8, by using a typical cudaStream to execute the subgraphs and interleaving them with computational operators, we enable communication within subgraphs to overlap with computation, thereby reducing latency.
6 Two-stage Sequence Assignment
@@
6.1 Sequence Packing within Pipeline
@@
Cost Model. For a given $P$, there is a maximum sequence length it can support, denoted as $\texttt{MaxLen}(P)$ (i.e., sequence exceeding this length will cause OOM on $P$), which can be determined in advance using our memory model (detailed in Appendix B.1). Additionally, we introduce a latency model, $\texttt{T}(l,P)$, which estimates the time taken to run a sequence of length $l$ on $P$ for a single forward and backward propagation of a single pipeline stage (detailed in Appendix B.2).
Solving for one $V$. The problem is easy to solve if the number of micro-batches $V$ is given. To be specific, as mentioned in §2.1, by employing the varlen functionality of FlashAttention v2 [18, 17], the latency of a packed sequence can be approximated as the sum of the latency of the individual original sequences. Meanwhile, to address the intra-pipeline imbalance, our goal is to minimize the duration of the longest-running micro-batch. Thus, we can formulate the following problem on top of the minimum sum partition (MSP) problem:
$\operatorname*{arg\,min}_{o_{i,j}}$ $\max\limits_{1\leq j\leq V}{\left(\sum\limits_{1\leq i\leq U}{o_{i,j}\times\texttt{T}\left(l_{i},P\right)}\right)\times(\texttt{PP}(P)-1+V)}$ (1)
@@
where $o_{i,j}\in\{0,1\}$ indicates whether the $i$-th sequence is packed into the $j$-th micro-batch. Compared to the MSP problem, the problem defined in Eq. (1) additionally ensures that the length of each packed sequence must not exceed $\texttt{MaxLen}(P)$, and we multiply the term $\texttt{PP}(P)-1+V$ to represent the propagation latency of the entire pipeline.
Enumeration of $V$. To locate the optimal number of micro-batches $V^{*}$, we can simply enumerate the value for $V\in[1,U]$, solve Eq. (1), and choose the best one. Note that Eq. (1) is an Integer Linear Programming (ILP) problem, but its complexity is only related to the number of sequences assigned to $P$ (i.e., $U$) and the number of micro-batches (i.e., $V$). These do not increase with cluster size, and the enumeration can be done in parallel, ensuring fast problem-solving even on large-scale clusters (detailed in Appendix E). Additionally, we can prune unnecessary enumerations, where a too-small $V$ results in a high bubble rate and a too-large $V$ leads to excessive kernel launch overheads. Detailed pruning methods are provided in Appendix C.
6.2 Sequence Dispatching across Pipelines
Next, we consider how the sequences in a mini-batch should be dispatched across the pipelines to address the inter-pipeline imbalance. Consider a specific strategy $S=P_{1}+\ldots+P_{D}$ satisfying $\texttt{MaxLen}\left(P_{1}\right)\geq\ldots\geq\texttt{MaxLen}\left(P_{D}\right)$. Assume a mini-batch contains $B$ sequences with lengths $\{l_{i}\}_{i=1}^{B}$. We denote $J_{i}$ as the number of pipelines that the $i$-th sequence can run on, satisfying $\texttt{MaxLen}\left(P_{J_{i}}\right)\geq l_{i}>\texttt{MaxLen}\left(P_{J_{i}+1}\right)$. Thus, $l_{i}$ can be dispatched to any of $P_{1},\ldots,P_{J_{i}}$. Since the pipelines execute concurrently, our goal is to dispatch the $B$ sequences across the $D$ pipelines while minimizing the propagation latency of the longest-running pipeline.
However, the propagation latency of each pipeline cannot be directly computed without solving the packing within it (i.e., §6.1). Fortunately, it is feasible to estimate the lower bound of latency. In particular, let $m_{i,j}\in\{0,1\}$ indicate whether the $i$-th sequence is assigned to $P_{j}$, then the lower bound of running time of the $j$-th pipeline can be expressed as (detailed derivation in Appendix D.1):
$\small\sum_{i=1}^{B}m_{i,j}\times\texttt{T}(l_{i},P_{j})+\texttt{T}\left(\max_{1\leq i\leq B}\left\{m_{i,j}\times l_{i}\right\},P_{j}\right)\times\left(\texttt{PP}\left(P_{j}\right)-1\right)$ (2)
@@ 6.2 Sequence Dispatching across Pipelines
$\operatorname*{arg\,min}_{m_{i,j}}$ $\max\limits_{1\leq j\leq D}\texttt{LowerBound}(\{m_{i,j}\}_{i=1}^{B},\{l_{i}\}_{i=1}^{B},P_{j})$ (3)
@@
The number of decision variables of Eq. (3) is linear w.r.t. number of sequences in a mini-batch (i.e., $B$) and the number of pipelines in the strategy (i.e., $D$). In practice, as the cluster scale expands, both will increase proportionally, making the assignment problem-solving cost unacceptable in large-scale scenarios. Fortunately, there are extensive studies about how to approximately solve assignment problems in polynomial time with a guaranteed error bound [24, 76, 12]. Motivated as such, we implement an approximate solver to Eq. (3), which utilizes multiple random trials to greedily dispatch sequences to pipelines and ultimately select the best trial. Empirical results demonstrate that the approximate solver achieves nearly the same objectives as directly solving Eq. (3), but with greater speed. More details about our approximate solver can be found in Appendix D and E.
7 Data Distribution-aware Strategy Proposal
@@
Figure 9: Dynamic programming with non-integer values. Different colors represent various $P$. The segments indicate sequences within each dataset interval, with their lengths corresponding to latency. Rounding the number of $P$ requires scaling the resulting latency.
Problem definition. To address both intra- and inter-iteration imbalances of data sampling, the proposal is based on two considerations. Firstly, since intra-iteration imbalance typically reflects the expected distribution of the dataset, we aim to preserve strategies that demonstrate an overall advantage over the entire dataset. Secondly, due to inter-iteration imbalance, characterized by variations in maximum sequence lengths, we aim to identify non-unique strategies across different maximum lengths. Then, we have the following problem.
Problem 1. Given the number of GPUs $N$, the dataset $\mathbb{D}$, and an arbitrary sequence length $L$. With sequence lengths randomly sampled within $[0,L]$ over $\mathbb{D}$, we seek the optimal heterogeneous strategy $S^{*}_{L|\mathbb{D}}$ from $\mathbb{S}$ that minimizes the expected cost. Then the proposed subset is defined as the union: $\hat{\mathbb{S}}=\bigcup_{L\leq L_{\text{max}}}\{S^{*}_{L|\mathbb{D}}\}$, where $L_{\text{max}}$ is the maximum sequence length to support (i.e., context length).
@@
Dynamic programming. We define $t[n][l]$ as the minimum achievable propagation latency using up to $n$ GPUs to process all sequences in the dataset with lengths not exceeding $l$. Our goal is to determine the value $t[N][L]$.
The initial states are obvious that $t[n][0]=0,\forall n>0$ and $t[0][l]=\infty,\forall l>0$. Next, we focus on the state transition for dynamic programming. Let $\mathbb{D}_{(l-l^{\prime},l]}$ denote the sub-dataset containing sequences within the interval $(l-l^{\prime},l]$, and assume the parallel scheme space comprises $K$ unique schemes $\{P_{k}\}_{k=1}^{K}$. Based on our simplification, we only need to consider assigning this sub-dataset to one kind of scheme. Suppose it is assigned to $d$ pipelines of $P_{k}$ and $P_{k}$ requires $\texttt{N}\left(P_{k}\right)$ GPUs, then this assignment consumes $d\times\texttt{N}\left(P_{k}\right)$ GPUs in total. Besides, we must ensure $\texttt{MaxLen}\left(P_{k}\right)\geq l$. Then the propagation latency of this sub-dataset can be approximated as $\frac{1}{d}\sum_{x\in\mathbb{D}_{(l-l^{\prime},l]}}\texttt{T}(x,P_{k})$. Meanwhile, the minimum propagation latency for the remaining sequences is $t\left[n-d\times\texttt{N}\left(P_{k}\right)\right]\left[l-l^{\prime}\right]$. Therefore, we have the following state transition expression:
$t[n][l]=\min\left(t[n-1][l],\min_{(k,d)\in{Cond}_{n,l},l^{\prime}<l}\left\{{T}_{k,d,l^{\prime}}\right\}\right),\ \textup{where}$ (4)
@@
Continuous relaxation. Based on the dynamic programming, we lift the restriction that each pipeline can only handle a sequence over a single length interval. Specifically, we execute the dynamic programming with non-integer values for $n$ and $d$. This means sequences in a length interval can use “a portion” of a pipeline, enabling a complete pipeline to support multiple intervals. Consequently, the final solution $S[N][L]$ might involve having non-integer numbers, as exemplified in Figure 9. In such cases, we round and adjust to include all nearby integer solutions in $\hat{\mathbb{S}}$, while ensuring that the total number of GPUs does not exceed $N$ [^4] The whole algorithm’s worst-case time complexity is $O(KN^{2}L_{\text{max}}^{2})$. In our experiments, we set the step size of $n$ and $d$ to 0.1 and $l$ to 128, with the actual time cost detailed in Appendix E.
@@
8 Evaluation
Figure 10: End-to-end evaluation. We test the workloads of three configurations, F1-F3, in each scenario. "Max Seq" represents the number of tokens in the longest sequence, while "Iter" denotes the total number of tokens sampled in a single iteration. Error bars indicate the standard deviation across iterations.
8.1 Experimental Setup
Environment. Our experiments are conducted with 8 GPU nodes, each containing 8 Nvidia A800 GPUs (64 GPUs in total). GPUs within a node are connected via NVLink with a bandwidth of 400 GB/s, and nodes are interconnected using Infini-Band with a bandwidth of 200 GB/s.
Baselines. We compare Hydraulis with Megatron-LM [77, 62, 46], DeepSpeed [72, 36], and HotSPa [28]. Both Megatron-LM and DeepSpeed employ a single, static homogeneous strategy throughout the training process. HotSPa uses heuristics to configure multiple homogeneous strategies and enables switching between them within every iteration. However, due to the overhead of switching, HotSPa is limited to setting 3-4 strategies. We tune the strategies for all competitors to achieve the best performance.
Models and datasets. Most existing large models are based on the decoder-only Transformer architecture. We select the widely used LLaMA2 series models and conduct experiments with 7B, 13B and 32B model sizes [79]. For datasets, we use CommonCrawl [78] and GitHub [16], which are widely used open-source datasets and have served as the major training corpora for many prestigious Transformer models.
Worloads and metrics. By default, we sample 100K tokens per iteration [^5] We consider 100K tokens per iteration for our 64-GPU cluster, scaled proportionally from the 16M used in LLaMA with an 8192-GPU cluster [25]. with a context length of 32K. And the Adam optimizer [45] is used. We conduct 20 warm-up iterations and then summarize the throughput (tokens per second) or per-iteration latency over 100 iterations.
Figure 11: Comparison of scalability across various factors, using the fastest one as the reference for relative changes. Shaded areas indicate the standard deviation across iterations.

### v2

@@ 5.3. Subgraphs
Although optimization-propagation disaggregation offers the ability to resolve the communication patterns needed for each heterogeneous strategy, it still poses the challenge of how to manage such diverse and complex optimization-propagation interactions, given that there would be numerous candidate heterogeneous strategies. To tackle this challenge, we introduce subgraphs, which are subsets of the complete computation graph. The key observation is that different strategies only have unique patterns in some specific operators, while most user-defined computational operators remain unchanged. Therefore, subgraphs are developed to encapsulate these unique operators from different strategies for efficient representation, management and scheduling.
Figure 10. The instantiation of Pull Parameter and Communicate Activation subgraphs during forward propagation. This involves inserting custom operators (e.g., send-receive, all-gather, LNK) to support arbitrary propagation strategies and optimization-propagation interactions.
@@
Subgraphs instantiation
As illustrated in Figure 10, users start by defining various computational operators to describe the training model. We then identify all potential points in the computational graph where different heterogeneous strategies might intervene and preserve several empty subgraphs (e.g., between transformer modules for potential pipeline parallel, before and after matrix multiplication for potential tensor parallel). Given a strategy $S$, the user-defined operators are directly integrated into a new, strategy-specific graph. We then determine how to instantiate all predefined subgraphs with appropriate operators. For example, as depicted at the top of Figure 10, a data transfer operator and a strategy-specific LNK operator are added to instantiate a Pull Parameter subgraph. This instantiation process ultimately forms a complete graph for a specific $S$, enabling different GPUs to extract and execute their respective operators.
Subgraphs scheduling
During execution, we optimize overlap by controlling the execution order of subgraphs and computational operators. Each subgraph can be treated as a cohesive unit and integrated into the scheduler as a fused operator. As illustrated in Figure 11, by using a typical cudaStream to execute the subgraphs and interleaving them with computational operators, we enable communication within subgraphs to overlap with computation, thereby reducing latency.
6. Two-stage Sequence Assignment
@@
Figure 11. Scheduling subgraphs with pipeline parallel (PP). In the first $P$ stage, Pull Parameter subgraphs are scheduled immediately after the last operators requiring the parameter, and Push Gradient subgraphs follow right after the operators that produce or accumulate the gradient. In subsequent $P$ stages, Pull Parameter and Push Gradient are grouped at the beginning and end, respectively, to utilize inherent $P$ bubbles. This scheduling scheme effectively hides communication latency within the Pull Parameter and Push Gradient subgraphs.
6.1. Sequence Packing within Pipeline
@@
Cost Model
Following prior works on automatic parallel training (Zheng et al., 2022; Li et al., 2025b; Lin et al., 2025; Zhao et al., 2025), we develop two profiling-based cost models to estimate memory usage and latency. For a given $P$, there is a maximum sequence length it can support, denoted as $\texttt{MaxLen}(P)$ (i.e., sequence exceeding this length will cause OOM on $P$), which can be determined in advance using our memory model (detailed in Appendix C.1 (hyd, 2025)). Additionally, we introduce a latency model, $\texttt{T}(l,P)$, which estimates the time taken to run a sequence of length $l$ on $P$ for a single forward and backward propagation of a single pipeline stage (detailed in Appendix C.2 (hyd, 2025)). Similar to prior works, re-profiling is required for different model architectures but is performed as a one-shot process per architecture. This overhead is worthwhile given the significant speedup in large-scale training. Moreover, this process is independent of the cluster size, as scaling the cluster only increases the number of pipelines, without affecting the set of valid $P$ or the sequence length $l$ used during training.
Solving for one $V$
The problem is easy to solve if the number of micro-batches $V$ is given. Specifically, as mentioned in §2.1, by employing the varlen functionality of FlashAttention v2 (Dao et al., 2022; Dao, 2023), the latency of a packed sequence can be approximated as the sum of latency for each original sequence. Meanwhile, to address the intra-pipeline imbalance, our goal is to minimize the duration of the longest-running micro-batch. Thus, we can formulate the following problem on top of the minimum sum partition (MSP) problem:
(1) $\operatorname*{arg\,min}_{o_{i,j}}$ $\max\limits_{1\leq j\leq V}{\left(\sum\limits_{1\leq i\leq U}{o_{i,j}\times\texttt{T}\left(l_{i},P\right)}\right)\times(\texttt{PP}(P)-1+V)}$
@@
where $o_{i,j}\in\{0,1\}$ indicates whether the $i$-th sequence is packed into the $j$-th micro-batch. Compared to the MSP problem, the problem defined in Eq. (1) additionally ensures that the length after packing must not exceed $\texttt{MaxLen}(P)$, and the term $\texttt{PP}(P)-1+V$ represents the propagation latency of the entire pipeline.
Enumeration of $V$
To locate the optimal number of micro-batches $V^{*}$, we can simply enumerate the value for $V\in[1,U]$, solve Eq. (1), and choose the best one.
Note that Eq. (1) is an Integer Linear Programming (ILP) problem, but its complexity is only related to the number of sequences assigned to $P$ (i.e., $U$) and the number of micro-batches (i.e., $V$). These do not increase with cluster size, and the enumeration can be done in parallel, ensuring fast problem-solving even on large-scale clusters (detailed in Appendix F (hyd, 2025)).
6.2. Sequence Dispatching across Pipelines
Next, we consider how the sequences in a mini-batch should be dispatched across the pipelines to address the inter-pipeline imbalance. A naïve approach is to bucket sequences by length. However, given the long-tail distribution of sequence lengths (where long sequences are rare, as shown in Figure 1), pipelines configured for long sequences may receive too few assignments during iterations dominated by short sequences. To improve load balancing, we need a more flexible approach that allows short sequences to be processed by any pipeline. Consider a specific strategy $S=P_{1}+\ldots+P_{D}$ satisfying $\texttt{MaxLen}\left(P_{1}\right)\geq\ldots\geq\texttt{MaxLen}\left(P_{D}\right)$. Assume a mini-batch contains $B$ sequences with lengths $\{l_{i}\}_{i=1}^{B}$. We denote $J_{i}$ as the number of pipelines that the $i$-th sequence can run on, satisfying $\texttt{MaxLen}\left(P_{J_{i}}\right)\geq l_{i}>\texttt{MaxLen}\left(P_{J_{i}+1}\right)$. Thus, $l_{i}$ can be dispatched to any of $P_{1},\ldots,P_{J_{i}}$. Since the pipelines execute concurrently, our goal is to dispatch the $B$ sequences across the $D$ pipelines while minimizing the propagation latency of the longest-running pipeline.
However, the propagation latency of each pipeline cannot be directly computed without solving the packing within it (i.e., §6.1). Fortunately, it is feasible to estimate the lower bound of latency. In particular, let $m_{i,j}\in\{0,1\}$ indicate whether the $i$-th sequence is assigned to $P_{j}$, then the lower bound of running time of the $j$-th pipeline can be expressed as (detailed derivation in Appendix E.1 (hyd, 2025)):
(2) $\small\sum_{i=1}^{B}m_{i,j}\times\texttt{T}(l_{i},P_{j})+\texttt{T}\left(\max_{1\leq i\leq B}\left\{m_{i,j}\times l_{i}\right\},P_{j}\right)\times\left(\texttt{PP}\left(P_{j}\right)-1\right)$
@@ 6.2. Sequence Dispatching across Pipelines
(3) $\operatorname*{arg\,min}_{m_{i,j}}$ $\max\limits_{1\leq j\leq D}\texttt{LowerBound}(\{m_{i,j}\}_{i=1}^{B},\{l_{i}\}_{i=1}^{B},P_{j})$
@@
The number of decision variables of Eq. (3) is linear w.r.t. number of sequences in a mini-batch (i.e., $B$) and the number of pipelines in the strategy (i.e., $D$). In practice, as the cluster scale expands, both will increase proportionally, making the assignment problem-solving cost unacceptable in large-scale scenarios. Fortunately, there are extensive studies about how to approximately solve assignment problems in polynomial time with a guaranteed error bound (Duan and Pettie, 2014; Shmoys and Tardos, 1993; Chen et al., 2020). Thus, we follow previous works to solve Eq. (3) approximately, which involves running multiple random trials to greedily dispatch sequences and selecting the best trial. Empirical results demonstrate that the approximation errors are consistently lower than 10% compared to directly solving Eq. (3). More details about our approximate solution can be found in Appendix E and Appendix F (hyd, 2025)
7. Data Distribution-aware Strategy Proposal
@@
Problem definition
Our proposal is based on two considerations regarding intra- and inter-iteration imbalances: First, since intra-iteration imbalance typically reflects the expected distribution of the dataset, we aim to preserve strategies that demonstrate an overall advantage over the entire dataset. Second, due to inter-iteration imbalance, characterized by variations in maximum sequence lengths, we aim to identify non-unique strategies across different maximum lengths. Then, we have the following problem.
Problem 1.
Given the number of GPUs $N$, the dataset $\mathbb{D}$, and an arbitrary sequence length $L$. With sequence lengths randomly sampled within $[0,L]$ over $\mathbb{D}$, we seek the optimal heterogeneous strategy $S^{*}_{L|\mathbb{D}}$ from $\mathbb{S}$ that minimizes the expected cost. Then the proposed subset is defined as the union: $\hat{\mathbb{S}}=\bigcup_{L\leq L_{\text{max}}}\{S^{*}_{L|\mathbb{D}}\}$, where $L_{\text{max}}$ is the maximum sequence length to support (i.e., context length).
@@
Dynamic programming
We define $t[n][l]$ as the minimum achievable propagation latency using up to $n$ GPUs to process all sequences in the dataset with lengths not exceeding $l$. Our goal is to determine the value $t[N][L]$. The initial states are obvious that $t[n][0]=0,\forall n>0$ and $t[0][l]=\infty,\forall l>0$. Next, we focus on the state transition for dynamic programming. Let $\mathbb{D}_{(l-l^{\prime},l]}$ denote the sub-dataset containing sequences within the interval $(l-l^{\prime},l]$, and assume the parallel scheme space comprises $K$ unique schemes $\{P_{k}\}_{k=1}^{K}$. Based on our simplification, we only need to consider assigning this sub-dataset to one kind of scheme. Suppose it is assigned to $d$ pipelines of $P_{k}$ and $P_{k}$ requires $\texttt{N}\left(P_{k}\right)$ GPUs, then this assignment consumes $d\times\texttt{N}\left(P_{k}\right)$ GPUs in total. Besides, we must ensure $\texttt{MaxLen}\left(P_{k}\right)\geq l$. Then the propagation latency of this sub-dataset can be approximated as $\frac{1}{d}\sum_{x\in\mathbb{D}_{(l-l^{\prime},l]}}\texttt{T}(x,P_{k})$. Meanwhile, the minimum propagation latency for the remaining sequences is $t\left[n-d\times\texttt{N}\left(P_{k}\right)\right]\left[l-l^{\prime}\right]$. Therefore, we have the following state transition expression:
(4) $t[n][l]=\min\left(t[n-1][l],\min_{(k,d)\in{Cond}_{n,l},l^{\prime}<l}\left\{{T}_{k,d,l^{\prime}}\right\}\right),\ \textup{where}$
@@
Figure 12. Dynamic programming with non-integer values. Different colors represent various $P$. The segments indicate sequences within each dataset interval, with their lengths corresponding to latency. Rounding the number of $P$ requires scaling the results.
Continuous relaxation
Based on the dynamic programming, we lift the restriction that each pipeline can only handle a sequence over a single length interval. Specifically, we execute the dynamic programming with non-integer values for $n$ and $d$. This means sequences in a length interval can use “a portion” of a pipeline, enabling a complete pipeline to support multiple intervals. Consequently, the final solution $S[N][L]$ might involve having non-integer numbers, as exemplified in Figure 12. In such cases, we round and adjust to include all nearby integer solutions in $\hat{\mathbb{S}}$, while ensuring that the total number of GPUs does not exceed NN. [^5] The whole algorithm’s worst-case time complexity is $O(KN^{2}L_{\text{max}}^{2})$. In our experiments, we set the step size of $n$ and $d$ to 0.1 and $l$ to 128, with the actual time cost detailed in Appendix F (hyd, 2025).
@@
Figure 13. End-to-end evaluation. We test the workloads of three configurations, F1-F3, in each scenario. “Max Seq” represents the number of tokens in the longest sequence, and “Iter” denotes the number of tokens sampled (a.k.a. batch size) in a single iteration. Error bars indicate the standard deviation across iterations.
8. Evaluation
Figure 14. Comparison of scalability across various factors, using the fastest one as the reference for relative changes. Shaded areas indicate the standard deviation across iterations.